OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = 4x – 5, what is f(x) ◦ g(x)? Answer a 18x + 3 b 8x – 7 c 8x – 15 d 8x + 7

hero (hero):

f(x)◦ g(x) = f(g(x)) which means everywhere you see x, put the entire g(x) in it: 2(4x-5) + 3 = 8x - 10 + 3 = 8x - 7